OpenStudy (anonymous):

Find the solutions of x^3-6x-5. Is this possible without graphing?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if you plug in 1 for function you should get 0, which means (x-1) is a factor

hero (hero):

Yes

OpenStudy (anonymous):

I say yes

OpenStudy (anonymous):

it's possible, just use wolfram or similar software. Or factor it yourself...

OpenStudy (anonymous):

sure it is if you want to read a really really really long formula for roots third degree polynomials.

OpenStudy (anonymous):

* -1 so it would be (x+1)

OpenStudy (anonymous):

but a good hint is if it was made up by a math teacher, try x = 1 and x = -1 first

OpenStudy (anonymous):

Well I started to type a reply and then seen all these people typing a reply at the same time lol

OpenStudy (anonymous):

my calculator says it has 3 roots, and 1 is 1 of them

OpenStudy (anonymous):

what imran said. try -1, say "oh wow i am shocked, it works!" then factor out (x+1)

OpenStudy (anonymous):

it would be uneven

hero (hero):

There's like twelve people here!

OpenStudy (anonymous):

That is the only rational root, could then use synthetic substitution to get the depressed equation of degree 2, then solve that

OpenStudy (anonymous):

(x+1)(x^2-x-5) is it factored

OpenStudy (anonymous):

\[x^3-6x-5=(x+1)(\text{something})\] and find the something. then quadratic formuala

OpenStudy (anonymous):

yeah do what mandolino said and you get the other 2 roots but they have \[\sqrt{reallyhardnumber - mumbojumbo}\] in them

OpenStudy (anonymous):

ya you could also use the quadratic formula...to assure if it is correct

OpenStudy (anonymous):

After synthetic with the root x=-1 you get x^2-x-5=0

OpenStudy (anonymous):

you can use synthetic division if you like, but you can also just think what it has to be

OpenStudy (anonymous):

Is there a better way to figure out that -1 is a root than guess-and-check?

OpenStudy (anonymous):

In this case, yes I agree

OpenStudy (anonymous):

x^2-x-5=0 has 2 real roots

OpenStudy (anonymous):

just like when you divide regular numbers. no there is not a better way than guess and check. only possible rational zeros are \[\pm1,\pm5\] so -1 is a good guess

OpenStudy (anonymous):

this has to be the most answered question i have ever seen!

OpenStudy (anonymous):

x = \[\frac{-(\sqrt{21}-1)}{2}\text{ or -1 or }\frac{\sqrt{21}+1}{2}\]

OpenStudy (anonymous):

x = \[\frac{-(\sqrt{21}-1)}{2}\text{ or -1 or }\frac{\sqrt{21}+1}{2}\]

OpenStudy (anonymous):

since when does the radical come first?

OpenStudy (anonymous):

my calculator prints it like that

OpenStudy (anonymous):

If you want to look it up in your book, satellite used what is called the rational roots (or zeros) theorem to get the list of possible rational roots.

OpenStudy (anonymous):

lol i thought it might be technology

OpenStudy (anonymous):

technology ftw

OpenStudy (anonymous):

x^3-6x-5=x^3-x-5x-5 x(x^2-1)-5(x+1) x(x-1)(x+1)-5(x+1) (x+1)(x(x-1)-5) (x+1)(x^2-x-5) x=-1 -b+-sqrt(b^2-4ac)/2a 1+-sqrt(1+20)/2 1+-sqrt(21)/2 x=-1 x=1+sqrt21/2 x=1-sqrt(21)/2